OpenStudy (anonymous):

can anyone help me with this ? Solve for x: 4 − (x + 2) < −3(x + 4) x < −7 x > −7 x < −9 x > −9

pooja195 (pooja195):

ok teat these like 2 sperate equations \[ -3(x+4)\] distribute

OpenStudy (anonymous):

so -3x-12 for that one correct?

OpenStudy (jdoe0001):

you'd solve \(\bf 4 − (x + 2) < −3(x + 4)\) same way you'd solve say \(\large \bf 4 − (x + 2) =0 −3(x + 4)\) more or less see what you get form the EQUAtion

pooja195 (pooja195):

where did u get the 0?

pooja195 (pooja195):

doesnt make a diffrence but lol whatever :P

OpenStudy (anonymous):

so -3x-12 for that one correct? @pooja195

OpenStudy (jdoe0001):

for the right-hand-side expansioin, that's correct, yes

OpenStudy (jdoe0001):

\(\bf 4 - (x + 2) < -3(x + 4)\implies 4-x-2<-3x-12\)

OpenStudy (anonymous):

I'm confused about the other side.

OpenStudy (jdoe0001):

-(x+2) => -1 * (x+2) -1 * x = -x -1 * +2 = -2 -(x+2) => -x-2

OpenStudy (anonymous):

I'm still confused about the answer

OpenStudy (jdoe0001):

well... how would you solve say 4 - x - 2 = -3x - 12 ?

OpenStudy (anonymous):

by first adding like terms right? @jdoe0001

OpenStudy (jdoe0001):

adding or subtracting, yes you'd want to leave the variable alone on either side alone by itself so you'd subtract and divide, as needed

OpenStudy (anonymous):

would i bring the "-x" over from the left side to add it to the "-3X" on the right side?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

and i would have 4-2<-2x-12 correct?

OpenStudy (jdoe0001):

yueap

OpenStudy (anonymous):

i was able to get the answer thank you very much @jdoe0001

OpenStudy (jdoe0001):

just keep in mind that with inequalities, when dividing, multiplying or exponentializing by a negative value you need to flip the inequality sign that's the only difference between an EQUAtion and an INEQUAlity

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf 4 - (x + 2) < -3(x + 4)\implies 4-x-2<-3x-12 \\ \quad \\ 4\cancel{-x{\color{brown}{ +x}}}-2<-3x{\color{brown}{ +x}}-12\implies 4-2<-2x-12 \\ \quad \\ 2{\color{brown}{ +12}}<-2x\cancel{-12{\color{brown}{ +12}}}\implies 14<-2x \\ \quad \\ {\color{brown}{ \textit{dividing by -2 on both sides, thus we }}}\ {\color{blue}{ flip }}\ the\ sign \\ \quad \\ \cfrac{14}{-2}{\color{blue}{ >}}\cfrac{\cancel{-2} x}{\cancel{-2}}\implies 7>x\)